### v1

@@ 4.2.2 Human perception rating
The set, and subset, of six indicators for urban perception (safe, lively, wealthy, beautiful, boring, and depressing) proposed in the PP2 dataset [6] have been widely and consistently adopted in related work on urban perception and human preferences [14, 11, 56, 57, 58, 43]. These traditional six perceptual indicators mainly capture people’s instantaneous and intuitive impressions of a place. Thus, we used them to facilitate comparative analysis with previous and future studies. However, we proposed and included four new indicators: live nearby, walk, cycle, and green. These additional indicators of living preference, cycling, walking, and greenery are presented as the following preference questions: “Which place looks like a place you want to live nearby?”, “Which place looks better to cycle?”, “Which place looks better to walk?”, and “Which place looks greener?”, to capture a proxy of belonging, bikeabiltiy, walkability, and subjective greenery, respectively. In contrast to the six traditional indicators, these four new ones reflect attitudes toward sustainability, livability, and active mobility, which are the central goals in many current urban studies [59, 60, 61]. Different studies tend to rely on these perceptual indicators as heuristics or proxies for walkability [60, 61] and bikeability [59]. Quantifying and assessing greenery remains crucial for planning and development, yet most studies rely on objective measurements such as Green View Index (GVI) despite its underestimation when compared to subjective measurements [62]. Our proposed indicators are meant to provide a more direct way of computing these dimensions that are commonly sought by the research community and practitioners. To support this, we hypothesize that these four indicators show a positive correlation with the positive indicators safe, lively, wealthy, and beautiful; and a negative correlation with the negative indicators boring and depressing.
@@ 4.4 Data analysis
When analyzing each demographic factor in isolation, i.e., computing Q scores based on only one demographic, we did not consider ratings from participants who answered “Prefer not to answer”(39 out of 1,000 participants in AHI), “Less than Secondary/High school” and “Other” (11 out of 1,000 participants in education level), and “A race/ethnicity not listed here” (69 out of 1,000 participants in the race and ethnicity question) due to their small sample size. On average, these excluded responses would have contributed only one additional rating to the Q score computation of each perceptual indicator for between 5 and 76 images per location. All subsequent analyses used all participants’ ratings.
@@
We created these demographic and personality subgroups for each country, and a set of subgroups by combining all countries too. While this nested subgrouping allowed for more detailed analysis, by reducing the number of available participants in each group, the number of pairwise comparisons per image per indicator was also reduced. Thus, after computing the Q scores in each nested subgroup, we retained all images with at least four pairwise comparisons, as this was the reported average comparison in PP2 [6]. We also used this number as a minimum number of samples, i.e., images, in each subgroup with more than four samples, i.e., rated images.
@@
Table 2: Annual Household Income (AHI) answers remapping.
@@
Second, we looked at how people living in the same place perceive different places. We compared the perception Q scores of imagery from multiple cities rated by participants from a single location, i.e., country. We also compared the predicted perception scores of all imagery by the ViT-PP2 perception model against the combined ratings from all participants without grouping them by location. We used the same statistical method from the previous analysis, Welch’s ANOVA, to find statistically significant differences, and we also used the same threshold of four minimum pairwise comparisons per image per indicator, and it resulted in at least 71% of the dataset, per country, usable. Third, in a similar vein, we compared how people from all five different countries perceived the same place. As this allowed us to use all participants’ ratings, we used a higher threshold of 22 pairwise comparisons. We used more than 51 images per city (63.75% of the available imagery per city). These perception Q scores were also compared with the ViT-PP2 predicted scores for imagery from each city. We performed these score comparisons with the output of a trained model as a way of comparing our collected ground truth against a one-size-fits-all prediction model. These comparisons were also based on Welch’s ANOVA statistical significance tests. Fourth, expanding on the second comparison of multi-city imagery and single-city participants, we looked into how people living in the same (or neighboring) city perceive other cities relative to their own. As this approach is on a similar vein to the second one, we used the same threshold and dataset size. We analysed the average change in perception of participants rating imagery from their own city compared to imagery from the other four cities. To debias the responses from cultural and location-based influences, we scaled the perception Q scores to z-scores among ratings for each city and indicator from all participants. This transformation avoids the cross-cultural perception bias as it looks at the scores within locations. We only included images with at least four pairwise comparisons per indicator, resulting in more than 50 images with Q scores per city per indicator (62.5% of the available SVI per city). Finally, we computed the Pearson correlation coefficient (R) between our proposed new four indicators and the existing six traditional ones to validate their usability and interpretability. For this, we calculated the perception Q scores using the ratings from all 1,000 participants and retained the images with at least 22 pairwise comparisons per indicator for more stable results as suggested by [34].
@@ Appendix E Extended results
Figure 10: Polynomial regression analyses between the green perceptual indicator and the six traditional ones: safe, lively, wealthy, beautiful, boring, and depressing. The best fitting curve between linear, quadratic, and cubic, in terms of $R^{2}$, is plotted in red. For linear models, $R$ values are also shown.

### v2

@@ 4.2.2 Human perception rating
The set, and subset, of six indicators for urban perception (safe, lively, wealthy, beautiful, boring, and depressing) proposed in the PP2 dataset [6] have been widely and consistently adopted in related work on urban perception and human preferences [14, 11, 57, 58, 59, 43]. These traditional six perceptual indicators mainly capture people’s instantaneous and intuitive impressions of a place. Thus, we used them to facilitate comparative analysis with previous and future studies. However, we proposed and included four new indicators: live nearby, walk, cycle, and green. These additional indicators of living preference, cycling, walking, and greenery are presented as the following preference questions: “Which place looks like a place you want to live nearby?”, “Which place looks better to cycle?”, “Which place looks better to walk?”, and “Which place looks greener?”, to capture a proxy of belonging, bikeabiltiy, walkability, and subjective greenery, respectively. In contrast to the six traditional indicators, these four new ones reflect attitudes toward sustainability, livability, and active mobility, which are the central goals in many current urban studies [60, 61, 62]. Different studies tend to rely on these perceptual indicators as heuristics or proxies for walkability [61, 62] and bikeability [60]. Moreover, quantifying and assessing greenery remains crucial for planning and development. The quality of green spaces has been linked to depression in older adults [63]; greenness in urban scenes was found to be a key indicator for assessing the urban renewal potential [64]; and green spaces offer critical insights for planning at the community level [65]. At the street level, most studies rely on objective measurements such as Green View Index (GVI), e.g., methods that asses pixel-based information, to asses the quality and quantity of green spaces. However, studies have found that GVI consistently underestimates greenery quantity when compared to subjective greenery evaluation [66] and subjective and visual measurements were found to contribute the most in the quality assessment of greenery [63, 64]. Moreover, [67] found that greenery could be perceived and valued differently by different genders and age groups, suggesting the need to investigate potential demographic differences in the perception of greenery in our study. Our proposed indicators are meant to provide a more direct way of computing these dimensions that are commonly sought by the research community and practitioners while maintaining human perspective which has been found to enable and support context-aware urban solutions [42, 11, 33]. To support this, we hypothesize that these four indicators show a positive correlation with the positive indicators safe, lively, wealthy, and beautiful; and a negative correlation with the negative indicators boring and depressing.
@@ 4.4 Data analysis
With this variable remapping, we also explored demographics interactions where possible. Using all five demographic factors (i.e., gender, age, income, education, race and ethnicity) to create nested subgroup reduces the available pairwise comparisons to a maximum of one pairwise comparison per image across all indicators and locations. Thus, we used only the combinations of ‘gender $\times$ age group $\times$ AHI’ as nested subgroups as these demographic factors are commonly explored in urban perception studies [13, 14, 15, 71] and among our demographics factors, they have fewer categories; two, four, and three, respectively. Nested subgroups based on these demographic factors have an average of six pairwise comparisons per image across all indicators and locations, above the threshold of four pairwise comparisons.
When analyzing each demographic factor in isolation and in interactions, i.e., computing Q scores based on a demographic subgroup, we did not consider ratings from participants who answered “Prefer not to answer”(39 out of 1,000 participants in AHI), “Less than Secondary/High school” and “Other” (11 out of 1,000 participants in education level), and “A race/ethnicity not listed here” (69 out of 1,000 participants in the race and ethnicity question) due to their small sample size. On average, these excluded responses would have contributed only one additional rating to the Q score computation of each perceptual indicator for between 5 and 76 images per location. All subsequent analyses used all participants’ ratings.
@@
We created these demographic and personality subgroups for each country, and a set of subgroups by combining all countries too. While this nested subgrouping allowed for more detailed analysis, by reducing the number of available participants in each group, the number of pairwise comparisons per image per indicator was also reduced. Thus, after computing the Q scores in each nested subgroup, we retained all images with at least four pairwise comparisons, as this was the reported average comparison in PP2 [6]. Since each demographic subgroup has a different number of samples, i.e., rated images with at least four pairwise comparison by participants with the same demographic characteristics, we prioritized a higher minimum number of samples, i.e., n $\geq 20$, where applicable but reported the minimum used in each demographic subgroup.
@@
Second, we looked at how people living in the same place perceive different places. We compared the perception Q scores of imagery from multiple cities rated by participants from a single location, i.e., country. We also compared the predicted perception scores of all imagery by the ViT-PP2 perception model against the combined ratings from all participants without grouping them by location. We used the same statistical method from the previous analysis, Welch’s ANOVA, to find statistically significant differences, and we also used the same threshold of four minimum pairwise comparisons per image per indicator, and it resulted in at least 71% of the dataset, per country, usable. Third, in a similar vein, we compared how people from all five different countries perceived the same place. As this allowed us to use all participants’ ratings, we used a higher threshold of 22 pairwise comparisons. We used more than 51 images per city (63.75% of the available imagery per city). These perception Q scores were also compared with the ViT-PP2 predicted scores for imagery from each city. We performed these score comparisons with the output of a trained model as a way of comparing our collected ground truth against a one-size-fits-all prediction model. These comparisons were also based on Welch’s ANOVA statistical significance tests. Fourth, expanding on the second comparison of multi-city imagery and single-city participants, we looked into how people living in the same (or neighboring) city perceive other cities relative to their own. As this approach is on a similar vein to the second one, we used the same threshold and dataset size. We analysed the average change in perception of participants rating imagery from their own city compared to imagery from the other four cities. To debias the responses from cultural and location-based influences, we scaled the perception Q scores to z-scores among ratings for each city and indicator from all participants. This transformation avoids the cross-cultural perception bias as it looks at the scores within locations. We only included images with at least four pairwise comparisons per indicator, resulting in more than 50 images with Q scores per city per indicator (62.5% of the available SVI per city). Finally, we performed linear, i.e., Pearson correlation coefficient (R), and non-linear, i.e., quadratic (normal and inverted U-shaped) and cubic, correlation analyses between our proposed new four indicators and the existing six traditional ones to validate their usability and interpretability. For this, we calculated the perception Q scores using the ratings from all 1,000 participants and retained the images with at least 22 pairwise comparisons per indicator for more stable results as suggested by [34].
@@ Appendix E Extended results
Figure 9: Statistical difference of perception scores by nested demographic. Perception Q scores are calculated from ratings by participants in each nested subgroup (combinations of ‘gender $\times$ age group $\times$ AHI’) for all (without location grouping) and each location. Locations with no significant differences in any indicator and nested demographic groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per subgroup) are not shown. Significance thresholds *$p<0.05$.
@@
(a) live nearby
(b) walk
(a) cycle
(b) green
Figure 12: Polynomial regression analyses between the proposed perceptual indicators ((a) live nearby, (b) walk, (c)cycle, and (d) green) and the six traditional ones: safe, lively, wealthy, beautiful, boring, and depressing. Best fitting curve between linear, quadratic (normal or inverted U-shaped) and cubic, in terms of $R^{2}$, are plotted in red. Linear fitting curve is shown alongside the best fitting model with $R$ values.